18 less than the product of twice a number and 3 is the same as 8 more than twice the difference of 11 and thrice the number. What is the number?
Answer by godeapurva(7)

Consider the number to be x
Given
LHS = RHS
We will form equation for LHS first.
18 less than twice product of a number and 3.
2X * 3 - 18 (LHS)
Now the RHS
8 more than twice the difference of 11 and thrice the number
2(11 - 3X) + 8 (RHS)
NOW
LHS=RHS
2x*3- 18 = 2(11- 3x) +8
6x-18 = 22- 6x + 8
12x = 48
x = 3